Shift report for evening 8/17/99

Sasha Glazov (glazov@hep.uchicago.edu)
Wed, 18 Aug 1999 00:03:53 -0500

Shift Summary for eve shift, 17/aug/99

On Shift: UE, SG.

Hours of beam available at all intensities: ~5h 50min
Hours of beam available at nominal intensity: ~5h 50min
Hours of E832 data (special runs not included): ~5h 40min
Hours of special runs (laser scans, muon runs...): 40min (laser scan)

Accounting of lost time: Tape change

========================================================================
Runs ended this shift:
Run 14253 -- Laser scan
Run 14256 -- 13.8M beam_e832

Tapes brought to FCC this shift:
(including tapes for online split)
none

========================================================================
Accesses during shift:

End of plan shutdown: work on CSi channels 216,226.

========================================================================
Overview of shift:
End of shutdown, Laser scan, E832 datataking

========================================================================
Description of Shift:

Fixing of CsI problems during shutdown time followed by laser scan.
Then E832 data taking with variable intensity (4 & 7 E12) and many
aborts.

run spill time comment
------------------------------------------------------------------------
14238 15:30 Shift starts -- no beam, down time continues
17:00 Problems with channels 216-226 are fixed.
The problem was due to loose cable conection.
17:08 Start laser scan
17:52 "Excellent!"
18:00 Put analysis magnet back on, initialize E832 run
18:07 Beam is back. But there are problems in crate 53,
right after initialization: UCLA scaler has yellow
light. Do manual init before the run start
14256 18:12 Kfrend, start data taking.
9-10 18:21 Beam is off target. Called MCR, they
say they are tuning. Also intensity is low (4E12).
22 Intensity goes to 7E12 but the rate of aborts
gets higher as well.
~160 21:55 MCR informs us that intensity will be low for a
while. (we notice decrease in the rate of aborts
as well)
22:22 MCR informs us that intensity will be high again
(we see also higher rate of aborts...)

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Monitor CsI channels 1137, 128.